trentin:

02.03] Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. (2 points) Equation 1 Equation 2 |5x + 6| = 41 |2x + 13| = 28

trentin:

@dude

dude:

Similar to what we just did with the past problems lets do equation 1 first \[|5x + 6| = 41\] We know that x can equal \[5x+6=-41\quad \quad \mathrm{or}\quad \:\quad \:5x+6=41\] For the first equation we can subtract 6 on both sides \[5x=-47\] Then we divide by 5 on both sides \[x=-\frac{47}{5}\]

dude:

For the second equation \[5x+6=41\] We also subtract 6 on both sides \[5x=35\] Then we divide by 5 on both sides \[x=7\]

dude:

Now for equation 2 \[|2x + 13| = 28]\] We know that x can equal \[2x+13=-28\quad \quad \mathrm{or}\quad \:\quad \:2x+13=28\] For the left equation \[2x+13=-28\] Subtract 13 on both sides \[2x=-41\] The we divide 2 on both sides \[x=-\frac{41}{2}\]

dude:

For the right equation \[2x+13=28\] Subtract 13 from both sides \[2x=15\] Divide 2 on both sides \[x=\frac{15}{2}\]

trentin:

Equation 1 and Equation 2 have the same number of solutions

dude:

So the solutions for equation #1 are \[x=−\frac{47}{5}, 7\] For equation #2 \[x=−\frac{41}{2}, \frac{15}{2}\]

dude:

Yes, they do have the same number of solutions

trentin:

yes

dude:

One difference would be that they do not intersect at the same points